In a random sample of 5 residents of the state of Tennessee, the mean waste recycled per person per day was 1.1 pounds with a standard deviation of 0.25 pounds. Determine the 99% confidence interval for the mean waste recycled per person per day for the population of Tennessee. Assume the population is approximately normal.

Step 1 of 2 :  

Find the critical value that should be used in constructing the confidence interval. Round your answer to three decimal places.

Step 2 of 2 :  

Construct the 99% confidence interval. Round your answer to one decimal place
Answer #1

Solution :

Given that,

Point estimate = sample mean = = 1.1

sample standard deviation = s = 0.25

sample size = n = 5

Degrees of freedom = df = n - 1 = 5 - 1 = 4

At 99% confidence level the t is ,

= 1 - 99% = 1 - 0.99 = 0.01

/ 2 = 0.01 / 2 = 0.005

t /2,df = t0.005,4 = 4.604

Margin of error = E = t/2,df * (s /n)

= 4.604 * (0.25 / 5)

= 0.5

The 99% confidence interval estimate of the population mean is,

- E < < + E

1.1 - 0.5 < < 1.1 + 0.5

0.6 < < 1.6

The 99% confidence interval : (0.6 , 1.6)